b. Given a population size of 4,3,2,4 and 2. Assume a sampling without replacement.

(a) find the population mean and standard deviation

(b) the sample mean and sample standard deviation.



1
Expert's answer
2021-05-19T17:26:15-0400

(a) Given population is 4,3,2,4,2


Population mean "\mu=\dfrac{4+3+2+4+2}{5}=\dfrac{15}{5}=3"


Standard deviation "\sigma=\sqrt{\dfrac{\sum (x-\mu)^2}{n}}"

"=\sqrt{\dfrac{(1)^2+0^2+(-1)^2+(1)^2+(-1)^2}{5}}\\[9pt]=\sqrt{\dfrac{4}{5}}\\[9pt]=\sqrt{0.8}=0.894"


(b) sample mean "x=\mu=3"


Sample standard deviation "s=\dfrac{\sigma}{\sqrt{n}}"


"=\dfrac{0.894}{\sqrt{5}}\\[9pt]=\dfrac{0.894}{2.245}=0.39"